A mixture of 9% disinfectant solution is to be made from 12% and 7% disinfectant solutions. How much of each solution should be used if 35 gallons of 9% solution are needed?
Answer by jorel1380(3719)

35x.09=3.15
p+q=35
.12p+.07q=3.15
12p+7q=315
7p+7q=245
5p=70
p=14
q=21
You need 14 gallons of 12% solution and 21 gallons of 7% solution..